# Mean Variance Efficient Collaborative Filtering for Stock Recommendations

### v2

@@ Appendix C Train and Validation Loss
$\text{BPR}_{\text{nov}}$ of Wasilewski et al. (2019) is a modification of BPR model (Rendle et al., 2009). BPR uses matrix factorization to predict ratings by minimizing $\mathop{\sum}\nolimits_{u,i,j\in{\mathcal{D}_{s}}}-\log\sigma\left({{{\hat{y}}_{ui}}-{{\hat{y}}_{uj}}}\right)$, where is the logistic sigmoid function, and $\mathcal{D}_{s}$} is a set of triples ($u,i,j)$ with ${y_{ui}}=1$ and ${y_{uj}}=0$. Wasilewski et al. (2019) defined a new set of triples $\mathcal{D}_{s}^{dist}=\{\left({u,i,j}\right):\left({u,i,j}\right)\in{\mathcal{D}_{s}}{\text{\;and\;dist}}\left({i,j}\right)<\tau\}$, where ${\text{dist}}\left({i,j}\right)$ is the distance between items $i$ and $j$, and $\tau$ is a distance threshold. They train the model with $\left({u,i,j}\right)$ sampled from $\mathcal{D}_{s}^{dist}$ with probability $\beta$ and sampled from $\mathcal{D}_{s}$} with probability $1-\beta$ to recommend items that are distinct from the items in the user’s preference history. For BPR and $\text{BPR}_{\text{nov}}$, we set $l=30$ and performed grid searches within the range $\lambda$$\in$0, 0.00001, 0.0001, 0.001 and ${\alpha}\in\left\{{0.001,0.005,0.01}\right\}$. The chosen values are ${\alpha}=0.005$ and ${\lambda}=0.00001$. We set the sampling probability $\beta$ to 80% for $\text{BPR}_{\text{nov}}$, and $\tau$ was set to 0.9 because it showed the best balance between novelty and recommendation performance among 0.8, 0.9, and 1.0. For graph based model, we used the same hyperparameters reported in paper and set the learning rate to 0.001, and the embedding size to 32. Figure 4 shows the train loss and the validation loss of $\text{MVECF}_{\text{wmf}}$ in one of the yearly sub-datasets (year 2015) of CRSP data. We can easily see that both models converge quite smoothly. It implies that restructuring MVECF into an ordinary WMF form makes enhancement in the computational efficiency. Similar results are found in all other datasets.

### v3

@@ 6. Conclusion
Acknowledgements.
This work was supported by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1I1A4069163).
@@ Appendix C Train and Validation Loss
$\text{BPR}_{\text{nov}}$ of Wasilewski et al. (Wasilewski2019) is a modification of BPR model (Rendle2009). BPR uses matrix factorization to predict ratings by minimizing $\mathop{\sum}\nolimits_{u,i,j\in{\mathcal{D}_{s}}}-\log\sigma\left({{{\hat{y}}_{ui}}-{{\hat{y}}_{uj}}}\right)$, where is the logistic sigmoid function, and $\mathcal{D}_{s}$} is a set of triples ($u,i,j)$ with ${y_{ui}}=1$ and ${y_{uj}}=0$. Wasilewski et al. (Wasilewski2019) defined a new set of triples $\mathcal{D}_{s}^{dist}=\{\left({u,i,j}\right):\left({u,i,j}\right)\in{\mathcal{D}_{s}}{\text{\;and\;dist}}\left({i,j}\right)<\tau\}$, where ${\text{dist}}\left({i,j}\right)$ is the distance between items $i$ and $j$, and $\tau$ is a distance threshold. They train the model with $\left({u,i,j}\right)$ sampled from $\mathcal{D}_{s}^{dist}$ with probability $\beta$ and sampled from $\mathcal{D}_{s}$} with probability $1-\beta$ to recommend items that are distinct from the items in the user’s preference history. For BPR and $\text{BPR}_{\text{nov}}$, we set $l=30$ and performed grid searches within the range $\lambda$$\in$0, 0.00001, 0.0001, 0.001 and ${\alpha}\in\left\{{0.001,0.005,0.01}\right\}$. The chosen values are ${\alpha}=0.005$ and ${\lambda}=0.00001$. We set the sampling probability $\beta$ to 80% for $\text{BPR}_{\text{nov}}$, and $\tau$ was set to 0.9 because it showed the best balance between novelty and recommendation performance among 0.8, 0.9, and 1.0. For graph based model, we used the same hyperparameters reported in paper and set the learning rate to 0.001, and the embedding size to 32. Figure 4 shows the train loss and the validation loss of $\text{MVECF}_{\text{wmf}}$ in one of the yearly sub-datasets (year 2015) of CRSP data. We can easily see that both models converge quite smoothly. It implies that restructuring MVECF into an ordinary WMF form makes enhancement in the computational efficiency. Similar results are found in all other datasets.